\begin{thebibliography}{}

\bibitem[\protect\citeauthoryear{Abdulkadiro{\u{g}}lu and
  S{\"o}nmez}{2003}]{AbSo03b}
Abdulkadiro{\u{g}}lu, A., and S{\"o}nmez, T.
\newblock 2003.
\newblock School choice: A mechanism design approach.
\newblock {\em American Economic Review} 93(3):729--747.

\bibitem[\protect\citeauthoryear{Abdulkadiro{\u{g}}lu, S{\"o}nmez, and
  {\"U}nver}{2004}]{ASU04a}
Abdulkadiro{\u{g}}lu, A.; S{\"o}nmez, T.; and {\"U}nver, M.~U.
\newblock 2004.
\newblock Room assignment-rent division: A market approach.
\newblock {\em Social Choice and Welfare} 22(3):515--538.

\bibitem[\protect\citeauthoryear{Airiau \bgroup et al\mbox.\egroup
  }{2019}]{AAC+19a}
Airiau, S.; Aziz, H.; Caragiannis, I.; Kruger, J.; Lang, J.; and Peters, D.
\newblock 2019.
\newblock Portioning using ordinal preferences: Fairness and efficiency.
\newblock In {\em Proceedings of the 28h International Joint Conference on
  Artificial Intelligence (IJCAI)}.

\bibitem[\protect\citeauthoryear{Alkan, Demange, and Gale}{1991}]{ADG91a}
Alkan, A.; Demange, G.; and Gale, D.
\newblock 1991.
\newblock Fair allocation of indivisible goods and criteria of justice.
\newblock {\em Econometrica} 59:1023--1039.

\bibitem[\protect\citeauthoryear{Amanatidis \bgroup et al\mbox.\egroup
  }{2015}]{AMNS15a}
Amanatidis, G.; Markakis, E.; Nikzad, A.; and Saberi, A.
\newblock 2015.
\newblock Approximation algorithms for computing maximin share allocations.
\newblock In {\em Proceedings of the 35th International Colloquium on Automata,
  Languages, and Programming (ICALP)},  39--51.

\bibitem[\protect\citeauthoryear{Atkinson}{1970}]{Atki70a}
Atkinson, A.
\newblock 1970.
\newblock On the measurement of inequality.
\newblock {\em J. of Economic Theory} 2(3):244--263.

\bibitem[\protect\citeauthoryear{Aziz and Klaus}{2019}]{AzKl19b}
Aziz, H., and Klaus, B.
\newblock 2019.
\newblock Random matching under priorities: Stability and no envy concepts.
\newblock {\em Social Choice and Welfare} 53(2):213--259.

\bibitem[\protect\citeauthoryear{Aziz and Lee}{2019}]{AzLe19a}
Aziz, H., and Lee, B.
\newblock 2019.
\newblock The expanding approvals rule: Improving proportional representation
  and monotonicity.
\newblock {\em Social Choice and Welfare}.

\bibitem[\protect\citeauthoryear{Aziz and Mackenzie}{2016}]{AzMa16c}
Aziz, H., and Mackenzie, S.
\newblock 2016.
\newblock A discrete and bounded envy-free cake cutting protocol for any number
  of agents.
\newblock In {\em Proceedings of the 57th Symposium on Foundations of Computer
  Science (FOCS)},  416--427.

\bibitem[\protect\citeauthoryear{Aziz and Rey}{2019}]{AzRe19a}
Aziz, H., and Rey, S.
\newblock 2019.
\newblock Almost group envy-free allocation of indivisible goods and chores.
\newblock {\em CoRR} abs/1907.09279.

\bibitem[\protect\citeauthoryear{Aziz \bgroup et al\mbox.\egroup
  }{2015}]{AGMW15a}
Aziz, H.; Gaspers, S.; Mackenzie, S.; and Walsh, T.
\newblock 2015.
\newblock Fair assignment of indivisible objects under ordinal preferences.
\newblock {\em Artificial Intelligence} 227:71--92.

\bibitem[\protect\citeauthoryear{Aziz \bgroup et al\mbox.\egroup
  }{2017a}]{ABC+16a}
Aziz, H.; Brill, M.; Conitzer, V.; Elkind, E.; Freeman, R.; and Walsh, T.
\newblock 2017a.
\newblock Justified representation in approval-based committee voting.
\newblock {\em Social Choice and Welfare} 48(2):461--485.

\bibitem[\protect\citeauthoryear{Aziz \bgroup et al\mbox.\egroup
  }{2017b}]{ARSW17a}
Aziz, H.; Rauchecker, G.; Schryen, G.; and Walsh, T.
\newblock 2017b.
\newblock Algorithms for max-min share fair allocation of indivisible chores.
\newblock In {\em Proceedings of the 31st AAAI Conference on Artificial
  Intelligence (AAAI)},  335--341.

\bibitem[\protect\citeauthoryear{Aziz \bgroup et al\mbox.\egroup
  }{2019a}]{ACI+18}
Aziz, H.; Caragiannis, I.; Igarashi, A.; and Walsh, T.
\newblock 2019a.
\newblock Fair allocation of indivisible goods and chores.
\newblock In {\em Proceedings of the 28th International Joint Conference on
  Artificial Intelligence (IJCAI)}.

\bibitem[\protect\citeauthoryear{Aziz \bgroup et al\mbox.\egroup
  }{2019b}]{AGSW19a}
Aziz, H.; Gaspers, S.; Sun, Z.; and Walsh, T.
\newblock 2019b.
\newblock From matching with diversity constraints to matching with regional
  quotas.
\newblock In {\em Proceedings of the 18th International Conference on
  Autonomous Agents and Multi-Agent Systems (AAMAS)}.

\bibitem[\protect\citeauthoryear{Aziz, Bogomolnaia, and Moulin}{2019}]{ABM19a}
Aziz, H.; Bogomolnaia, A.; and Moulin, H.
\newblock 2019.
\newblock Fair mixing: the case of dichotomous preferences.
\newblock In {\em Proceedings of the 20th ACM Conference on Economics and
  Computation (ACM-EC)},  753--781.

\bibitem[\protect\citeauthoryear{Aziz, Lee, and Talmon}{2018}]{ALT18a}
Aziz, H.; Lee, B.~E.; and Talmon, N.
\newblock 2018.
\newblock Proportionally representative participatory budgeting: Axioms and
  algorithms.
\newblock In {\em Proceedings of the 17th International Conference on
  Autonomous Agents and MultiAgent Systems, {AAMAS} 2018, Stockholm, Sweden,
  July 10-15, 2018},  23--31.

\bibitem[\protect\citeauthoryear{Aziz}{2016}]{Aziz16a}
Aziz, H.
\newblock 2016.
\newblock Computational social choice: Some current and new directions.
\newblock In {\em Proceedings of the 25th International Joint Conference on
  Artificial Intelligence (IJCAI)},  4054--4057.

\bibitem[\protect\citeauthoryear{Aziz}{2019a}]{Aziz18a}
Aziz, H.
\newblock 2019a.
\newblock A probabilistic approach to voting, allocation, matching, and
  coalition formation.
\newblock In Laslier, J.-F.; Moulin, H.; Sanver, R.; and Zwicker, W.~S., eds.,
  {\em The Future of Economic Design}. Springer-Verlag.
\newblock Forthcoming.

\bibitem[\protect\citeauthoryear{Aziz}{2019b}]{Aziz19b}
Aziz, H.
\newblock 2019b.
\newblock A rule for committee selection with soft diversity constraints.
\newblock {\em Group Decision and Negotiation}  1--8.

\bibitem[\protect\citeauthoryear{Balinski and Young}{1982}]{BaYo82a}
Balinski, M., and Young, H.~P.
\newblock 1982.
\newblock {\em Fair Representation: {M}eeting the Ideal of One Man, One Vote}.
\newblock Yale University Press.
\newblock (2nd Edition [with identical pagination], Brookings Institution
  Press, 2001).

\bibitem[\protect\citeauthoryear{Barman and Krishnamurthy}{2017}]{BaKr17}
Barman, S., and Krishnamurthy, S.~K.
\newblock 2017.
\newblock Approximation algorithms for maximin fair division.
\newblock In {\em Proceedings of the 18th ACM Conference on Economics and
  Computation (ACM-EC)}.

\bibitem[\protect\citeauthoryear{Barman, Krishnamurthy, and
  Vaish}{2018}]{BKV18a}
Barman, S.; Krishnamurthy, S.~K.; and Vaish, R.
\newblock 2018.
\newblock Finding fair and efficient allocations.
\newblock In {\em Proceedings of the 2018 {ACM} Conference on Economics and
  Computation, Ithaca, NY, USA, June 18-22, 2018},  557--574.

\bibitem[\protect\citeauthoryear{Benabbou \bgroup et al\mbox.\egroup
  }{2019}]{BCEZ19}
Benabbou, N.; Chakraborty, M.; Elkind, E.; and Zick, Y.
\newblock 2019.
\newblock Fairness towards groups of agents in the allocation of indivisible
  items.
\newblock In {\em Proceedings of the 28th International Joint Conference on
  Artificial Intelligence (IJCAI)}.

\bibitem[\protect\citeauthoryear{Biro \bgroup et al\mbox.\egroup
  }{2010}]{BFIM10}
Biro, P.; Fleiner, T.; Irving, R.~W.; and Manlove, D.~F.
\newblock 2010.
\newblock The college admissions problem with lower and common quotas.
\newblock {\em Theoretical Computer Science} 411(34--36):3136--3153.

\bibitem[\protect\citeauthoryear{Bogomolnaia \bgroup et al\mbox.\egroup
  }{2017}]{BMSY17}
Bogomolnaia, A.; Moulin, H.; Sandomirskiy, F.; and Yanovskaya, E.
\newblock 2017.
\newblock Competitive division of a mixed manna.
\newblock {\em Econometrica} 85(6):1847--1871.

\bibitem[\protect\citeauthoryear{Bouveret \bgroup et al\mbox.\egroup
  }{2017}]{BCE+17a}
Bouveret, S.; Cechl{\'{a}}rov{\'{a}}, K.; Elkind, E.; Igarashi, A.; and Peters,
  D.
\newblock 2017.
\newblock Fair division of a graph.
\newblock In {\em Proceedings of the Twenty-Sixth International Joint
  Conference on Artificial Intelligence, {I ]JCAI} 2017, Melbourne, Australia,
  August 19-25, 2017},  135--141.

\bibitem[\protect\citeauthoryear{Bouveret, Chevaleyre, and
  Maudet}{2016}]{BCM15a}
Bouveret, S.; Chevaleyre, Y.; and Maudet, N.
\newblock 2016.
\newblock Fair allocation of indivisible goods.
\newblock In Brandt, F.; Conitzer, V.; Endriss, U.; Lang, J.; and Procaccia,
  A.~D., eds., {\em Handbook of Computational Social Choice}. Cambridge
  University Press.
\newblock chapter~12.

\bibitem[\protect\citeauthoryear{Brams and Togman}{1996}]{BrTo96a}
Brams, S.~J., and Togman, J.~M.
\newblock 1996.
\newblock Camp david: Was the agreement fair?
\newblock {\em Conflict Management and Peace Science} 15(1):99--112.

\bibitem[\protect\citeauthoryear{Brams}{2008}]{Bram08a}
Brams, S.~J.
\newblock 2008.
\newblock {\em Mathematics and Democracy: Designing Better Voting and
  Fair-Division Procedures}.
\newblock Princeton University Press.

\bibitem[\protect\citeauthoryear{Brandt}{2017}]{Bran17a}
Brandt, F.
\newblock 2017.
\newblock Rolling the dice: {R}ecent results in probabilistic social choice.
\newblock In Endriss, U., ed., {\em Trends in Computational Social Choice}. AI
  Access.
\newblock chapter~1,  3--26.

\bibitem[\protect\citeauthoryear{Budish}{2011}]{Budi11a}
Budish, E.
\newblock 2011.
\newblock The combinatorial assignment problem: Approximate competitive
  equilibrium from equal incomes.
\newblock {\em Journal of Political Economy} 119(6):1061--1103.

\bibitem[\protect\citeauthoryear{Caragiannis \bgroup et al\mbox.\egroup
  }{2016}]{CKM+16a}
Caragiannis, I.; Kurokawa, D.; Moulin, H.; Procaccia, A.~D.; Shah, N.; and
  Wang, J.
\newblock 2016.
\newblock {The Unreasonable Fairness of Maximum Nash Welfare}.
\newblock In {\em Proceedings of the 17th ACM Conference on Economics and
  Computation (ACM-EC)},  305--322.

\bibitem[\protect\citeauthoryear{Cole and Gkatzelis}{2015}]{CoGz15a}
Cole, R., and Gkatzelis, V.
\newblock 2015.
\newblock Approximating the {N}ash social welfare with indivisible items.
\newblock In {\em Proceedings of the 47th Annual ACM Symposium on Theory of
  Computing (STOC)},  371--380.
\newblock ACM Press.

\bibitem[\protect\citeauthoryear{Conitzer \bgroup et al\mbox.\egroup
  }{2019}]{CFS+19}
Conitzer, V.; Freeman, R.; Shah, N.; and Vaughan, J.~W.
\newblock 2019.
\newblock Group fairness for the allocation of indivisible goods.
\newblock In {\em Proceedings of the 33rd AAAI Conference on Artificial
  Intelligence (AAAI)}.

\bibitem[\protect\citeauthoryear{Conitzer, Freeman, and Shah}{2017}]{CF017}
Conitzer, V.; Freeman, R.; and Shah, N.
\newblock 2017.
\newblock Fair public decision making.
\newblock In {\em Proceedings of the 2017 {ACM} Conference on Economics and
  Computation, {EC} '17},  629--646.

\bibitem[\protect\citeauthoryear{Dehghani \bgroup et al\mbox.\egroup
  }{2018}]{DFHY18a}
Dehghani, S.; Farhadi, A.; Hajiaghayi, M.~T.; and Yami, H.
\newblock 2018.
\newblock Envy-free chore division for an arbitrary number of agents.
\newblock In {\em Proceedings of the Twenty-Ninth Annual {ACM-SIAM} Symposium
  on Discrete Algorithms, {SODA} 2018, New Orleans, LA, USA, January 7-10,
  2018},  2564--2583.

\bibitem[\protect\citeauthoryear{Ehlers \bgroup et al\mbox.\egroup
  }{2014}]{EHYY14a}
Ehlers, L.; Hafalir, I.~E.; Yenmezb, M.~B.; and Yildirimc, M.~A.
\newblock 2014.
\newblock School choice with controlled choice constraints: Hard bounds versus
  soft bounds.
\newblock {\em Journal of Economic Theory} 153:648---683.

\bibitem[\protect\citeauthoryear{Fain, Goel, and Munagala}{2016}]{FGM16a}
Fain, B.; Goel, A.; and Munagala, K.
\newblock 2016.
\newblock The core of the participatory budgeting problem.
\newblock In {\em Web and Internet Economics - 12th International Conference,
  {WINE} 2016, Montreal, Canada, December 11-14, 2016, Proceedings},  384--399.

\bibitem[\protect\citeauthoryear{Faliszewski \bgroup et al\mbox.\egroup
  }{2017}]{FSST17a}
Faliszewski, P.; Skowron, P.; Slinko, A.; and Talmon, N.
\newblock 2017.
\newblock Multiwinner voting: A new challenge for social choice theory.
\newblock In Endriss, U., ed., {\em Trends in Computational Social Choice}.
\newblock chapter~2.

\bibitem[\protect\citeauthoryear{Fluschnik \bgroup et al\mbox.\egroup
  }{2017}]{FST+17a}
Fluschnik, T.; Skowron, P.; Triphaus, M.; and Wilker, K.
\newblock 2017.
\newblock Fair knapsack.
\newblock {\em CoRR} abs/1711.04520.

\bibitem[\protect\citeauthoryear{Gal \bgroup et al\mbox.\egroup
  }{2017}]{GMPZ17a}
Gal, Y.; Mash, M.; Procaccia, A.~D.; and Zick, Y.
\newblock 2017.
\newblock Which is the fairest (rent division) of them all?
\newblock {\em J. {ACM}} 64(6):39:1--39:22.

\bibitem[\protect\citeauthoryear{Ghodsi \bgroup et al\mbox.\egroup
  }{2011}]{GZH+11a}
Ghodsi, A.; Zaharia, M.; Hindman, B.; Konwinski, A.; Shenker, S.; and Stoica,
  I.
\newblock 2011.
\newblock Dominant resource fairness: Fair allocation of multiple resource
  types.
\newblock In {\em Proceedings of the 8th USENIX Symposium on Networked Systems
  Design and Implementation}.

\bibitem[\protect\citeauthoryear{Ghodsi \bgroup et al\mbox.\egroup
  }{2018}]{SGHSY18}
Ghodsi, M.; HajiAghayi, M.; Seddighin, M.; Seddighin, S.; and Yami, H.
\newblock 2018.
\newblock Fair allocation of indivisible goods: Improvements and
  generalizations.
\newblock In {\em Proceedings of the 19th ACM Conference on Economics and
  Computation (ACM-EC)}.
\newblock ACM Press.

\bibitem[\protect\citeauthoryear{Goto \bgroup et al\mbox.\egroup
  }{2016}]{GIKY+16a}
Goto, M.; Iwasaki, A.; Kawasaki, Y.; Kurata, R.; Yasuda, Y.; and Yokoo, M.
\newblock 2016.
\newblock Strategyproof matching with regional minimum and maximum quotas.
\newblock {\em Artificial intelligence} 235:40--57.

\bibitem[\protect\citeauthoryear{Huang and Lu}{2019}]{XiLu19a}
Huang, X., and Lu, P.
\newblock 2019.
\newblock An algorithmic framework for approximating maximin share allocation
  of chores.
\newblock {\em CoRR} abs/1907.04505.

\bibitem[\protect\citeauthoryear{Kamada and Kojima}{2015}]{KaKo15}
Kamada, Y., and Kojima, F.
\newblock 2015.
\newblock Efficient matching under distributional constraints: Theory and
  applications.
\newblock {\em American Economic Review} 105(1):67--99.

\bibitem[\protect\citeauthoryear{Kesten and Unver}{2015}]{KeUn15a}
Kesten, O., and Unver, U.
\newblock 2015.
\newblock A theory of school choice lotteries.
\newblock {\em Theoretical Economics}  543---595.

\bibitem[\protect\citeauthoryear{Kleinberg \bgroup et al\mbox.\egroup
  }{2018}]{KLR18a}
Kleinberg, J.; Ludwig, J.; Mullainathan, S.; and Rambachan, A.
\newblock 2018.
\newblock Algorithmic fairness.
\newblock {\em American Economic Association Papers and Proceedings}
  108:22--27.

\bibitem[\protect\citeauthoryear{Kleinberg, Rabani, and Tardos}{2001}]{KRT01a}
Kleinberg, J.~M.; Rabani, Y.; and Tardos, {\'{E}}.
\newblock 2001.
\newblock Fairness in routing and load balancing.
\newblock {\em J. Comput. Syst. Sci.} 63(1):2--20.

\bibitem[\protect\citeauthoryear{Kurata \bgroup et al\mbox.\egroup
  }{2015}]{KHIY15a}
Kurata, R.; Hamada, N.; Iwasaki, A.; and Yokoo, M.
\newblock 2015.
\newblock Controlled school choice with soft bounds and overlapping types.
\newblock In {\em Proceedings of the 29th AAAI Conference on Artificial
  Intelligence (AAAI)},  951--957.

\bibitem[\protect\citeauthoryear{Lang and Rothe}{2016}]{LaRo16}
Lang, J., and Rothe, J.
\newblock 2016.
\newblock Fair division of indivisible goods.
\newblock In Rothe, J., ed., {\em Economics and Computation}.
\newblock  493--550.

\bibitem[\protect\citeauthoryear{Manlove}{2013}]{Manl13a}
Manlove, D.~F.
\newblock 2013.
\newblock {\em Algorithmics of Matching Under Preferences}.
\newblock World Scientific Publishing Company.

\bibitem[\protect\citeauthoryear{Moulin}{2003}]{Moul03a}
Moulin, H.
\newblock 2003.
\newblock {\em Fair Division and Collective Welfare}.
\newblock The MIT Press.

\bibitem[\protect\citeauthoryear{Moulin}{2019}]{Moul18}
Moulin, H.
\newblock 2019.
\newblock Fair division in the age of internet.
\newblock {\em Annual Review of Economics} 11:407--441.

\bibitem[\protect\citeauthoryear{Nguyen \bgroup et al\mbox.\egroup
  }{2014}]{NNRJ14a}
Nguyen, N.-T.; Nguyen, T.-T.; Roos, M.; and Rothe, J.
\newblock 2014.
\newblock Computational complexity and approximability of social welfare
  optimization in multiagent resource allocation.
\newblock {\em Autonomous Agents and Multi-Agent Systems} 28(2):256--289.

\bibitem[\protect\citeauthoryear{Oh, Procaccia, and Suksompong}{2019}]{OPS19a}
Oh, H.; Procaccia, A.~D.; and Suksompong, W.
\newblock 2019.
\newblock Fairly allocating many goods with few queries.
\newblock In {\em Proceedings of the 33rd AAAI Conference on Artificial
  Intelligence (AAAI)},  2141--2148.

\bibitem[\protect\citeauthoryear{Parkes, Procaccia, and Shah}{2013}]{PPS12a}
Parkes, D.~C.; Procaccia, A.~D.; and Shah, N.
\newblock 2013.
\newblock Beyond dominant resource fairness: Extensions, limitations, and
  indivisibilities.
\newblock In {\em Proceedings of the 13th ACM Conference on Electronic Commerce
  (ACM-EC)},  808--825.

\bibitem[\protect\citeauthoryear{Plaut and Roughgarden}{2018}]{PlRo18}
Plaut, B., and Roughgarden, T.
\newblock 2018.
\newblock Almost envy-freeness with general valuations.
\newblock In {\em Proceedings of the 29th Annual ACM-SIAM Symposium on Discrete
  Algorithms (SODA)}.

\bibitem[\protect\citeauthoryear{Procaccia and Wang}{2014}]{PrWa14a}
Procaccia, A.~D., and Wang, J.
\newblock 2014.
\newblock Fair enough: Guaranteeing approximate maximin shares.
\newblock In {\em Proceedings of the 15th ACM Conference on Economics and
  Computation (ACM-EC)},  675--692.
\newblock ACM Press.

\bibitem[\protect\citeauthoryear{Procaccia}{2013}]{Proc12a}
Procaccia, A.~D.
\newblock 2013.
\newblock Cake cutting: Not just child's play.
\newblock {\em Communications of the ACM} 56(7):78--87.

\bibitem[\protect\citeauthoryear{Procaccia}{2016}]{Proc15a}
Procaccia, A.~D.
\newblock 2016.
\newblock Cake cutting algorithms.
\newblock In Brandt, F.; Conitzer, V.; Endriss, U.; Lang, J.; and Procaccia,
  A.~D., eds., {\em Handbook of Computational Social Choice}. Cambridge
  University Press.
\newblock chapter~13.

\bibitem[\protect\citeauthoryear{Rawls}{1971}]{Rawl71a}
Rawls, J.
\newblock 1971.
\newblock {\em A Theory of Justice}.
\newblock Harvard University Press.

\bibitem[\protect\citeauthoryear{Rescher}{2002}]{Resc02a}
Rescher, N.
\newblock 2002.
\newblock {\em Fairness: Theory and Practice of Distributive Justice}.
\newblock Transaction Publishers.

\bibitem[\protect\citeauthoryear{Robertson and Webb}{1998}]{RoWe98a}
Robertson, J.~M., and Webb, W.~A.
\newblock 1998.
\newblock {\em Cake Cutting Algorithms: Be Fair If You Can}.
\newblock A. K. Peters.

\bibitem[\protect\citeauthoryear{Roth}{2008}]{Roth08a}
Roth, A.~E.
\newblock 2008.
\newblock Deferred acceptance algorithms: history, theory, practice, and open
  questions.
\newblock {\em International Journal of Game Theory} 36:537---569.

\bibitem[\protect\citeauthoryear{Sandomirskiy and Segal-Halevi}{2019}]{SaSe19b}
Sandomirskiy, F., and Segal-Halevi, E.
\newblock 2019.
\newblock Fair division with minimal sharing.
\newblock {\em CoRR} abs/1908.01669.

\bibitem[\protect\citeauthoryear{Segal-Halevi and Suksompong}{2019}]{SeSu19a}
Segal-Halevi, E., and Suksompong, W.
\newblock 2019.
\newblock Democratic fair allocation of indivisible goods.
\newblock {\em Artificial Intelligence} 277:103--167.

\bibitem[\protect\citeauthoryear{Segal{-}Halevi, Hassidim, and
  Aumann}{2015}]{SHA15a}
Segal{-}Halevi, E.; Hassidim, A.; and Aumann, Y.
\newblock 2015.
\newblock Envy-free cake-cutting in two dimensions.
\newblock In {\em Proceedings of the 29th AAAI Conference on Artificial
  Intelligence (AAAI)},  1021--1028.

\bibitem[\protect\citeauthoryear{Shin and Park}{2019}]{ShPa19a}
Shin, D., and Park, Y.~J.
\newblock 2019.
\newblock Role of fairness, accountability, and transparency in algorithmic
  affordance.
\newblock {\em Computers in Human Behavior} 98:277--284.

\bibitem[\protect\citeauthoryear{Thomson}{2011}]{Thom11b}
Thomson, W.
\newblock 2011.
\newblock Chapter twenty-one - fair allocation rules.
\newblock In Arrow, K.~J.; Sen, A.; and Suzumura, K., eds., {\em Handbook of
  Social Choice and Welfare}, volume~2 of {\em Handbook of Social Choice and
  Welfare}. Elsevier.
\newblock  393 -- 506.

\bibitem[\protect\citeauthoryear{Thomson}{2016}]{Thom15a}
Thomson, W.
\newblock 2016.
\newblock Introduction to the theory of fair allocation.
\newblock In Brandt, F.; Conitzer, V.; Endriss, U.; Lang, J.; and Procaccia,
  A.~D., eds., {\em Handbook of Computational Social Choice}. Cambridge
  University Press.
\newblock chapter~11.

\bibitem[\protect\citeauthoryear{Varian}{1974}]{Vari74a}
Varian, H.~R.
\newblock 1974.
\newblock Equity, envy, and efficiency.
\newblock {\em Journal of Economic Theory} 9:63--91.

\end{thebibliography}

\def\year{2020}\relax
\documentclass[letterpaper]{article} % DO NOT CHANGE THIS
\usepackage{aaai20-report}  % DO NOT CHANGE THIS
\usepackage{times}  % DO NOT CHANGE THIS
\usepackage{helvet} % DO NOT CHANGE THIS
\usepackage{courier}  % DO NOT CHANGE THIS
\usepackage[hyphens]{url}  % DO NOT CHANGE THIS
\usepackage{graphicx} % DO NOT CHANGE THIS
\usepackage{graphicx}  % DO NOT CHANGE THIS
\usepackage{helvet}  %Required
\usepackage{courier}  %Required
\usepackage[hyphens]{url}
\usepackage{graphicx}  %Required
\usepackage{nicefrac}
\usepackage{booktabs}

\usepackage{algorithm,algorithmic}

\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{mathtools}
\usepackage{amsfonts}
\usepackage{multicol}

		\newcommand{\citet}[1]{\citeauthor{#1}~\shortcite{#1}}
		\newcommand{\citep}{\cite}
\newcommand{\myOmit}[1]{}

\usepackage{color}

\title{Developments in Multi-Agent Fair Allocation}
\author{Haris Aziz\\UNSW Sydney and Data61\\
    Sydney, Australia\\
       haziz@cse.unsw.edu.au}
\begin{document}

\maketitle

\begin{abstract}
Fairness is becoming an increasingly important concern when designing markets, allocation procedures, and computer systems. I survey some recent developments in the field of multi-agent fair allocation. 
\end{abstract}

\section{Introduction}

Many important decisions in our lives are increasingly being made by computers. These decisions include whether we get a certain job or loan. It becomes imperative then that computer systems provide guarantees that the decision has been made transparently and fairly. For many settings such as two-sided matching, fairness considerations based on justified envy or diversity guarantees are enshrined in the letter of the law. Even for smaller day to day decisions, 
the appetite for fairness guarantees is naturally present among people.\footnote{See, for example, the interest in fair allocation portal \url{www.spliddit.org/}).} 
At a bigger scale, the Internet leads to peer to peer interactions where fairness comes up as an important concern~(see, e.g. \citet{Moul18} and \citet{KRT01a}).

The issue of what is fair and how to achieve fairness has been studied in several disciplines including economics~\citep{Moul03a,Thom11b}, political science~\citep{Bram08a,BaYo82a} and philosophy~\citep{Resc02a,Rawl71a}.\footnote{Despite the richness of fairness concepts and principles in these literatures, I will only focus on a handful of fairness concepts in this paper. Other fairness ideas include indices to measure inequality (see, e.g. \citet{Atki70a}) or market-based approaches in which agents have equal budgets (see, e.g. \citet{Vari74a}). } 

I mention some recent developments in fair allocation in multi-agent systems. In particular, I discuss allocation of indivisible items (Section~\ref{sec:indiv}), allocation of divisible items (Section~\ref{sec:div}), two-sided matching (Section~\ref{sec:match}), and fairness in other social choice settings (Section~\ref{sec:sc}). Finally, I wrap up the discussion in Section~\ref{sec:discuss}.

\section{Allocation of Indivisible Items}
\label{sec:indiv}

When indivisible items are allocated among agents, guaranteeing the existence of fair allocations is impossible for many desirable notions of fairness. For example, if there are two agents and one indivisible valued good, allocating the good to either of the two agents will make the other agent envious. In other words, an envy-free allocation may not exist. Another challenge is computational. Suppose agents have positive additive utilities over the items. Then the problem
of checking whether there exists an envy-free allocation is NP-complete even for 1-0 utilities~\citep{AGMW15a}. Similarly the problems of computing allocations that maximize the egalitarian welfare~(utility of the worst off agent) or Nash welfare (geometric mean of the utilities of agents) is NP-hard~(see, e.g. \citet{NNRJ14a}). 
%Does there exists an allocation in which each agent gets at least a certain amount of utility? 

Given these challenges, a couple of influential fairness concepts were proposed by \citet{Budi11a}. The first one is called \emph{maxmin share (MMS)} fairness. The \emph{maxmin fair share} of an agent is the best she can guarantee for herself if she is allowed to partition the items into bundles for the agents but then receives her least preferred bundle. An allocation satisfies MMS fairness if each agent's allocation is at least as preferred to her as her maxmin fair share. Although an MMS fair allocation is not guaranteed to exist even for additive utilities, the concept gives rise to approximation versions in which each agent aspires to get a certain percentage of her maxmin fair share~\citep{PrWa14a}. This has led to several results including new algorithms that are faster, simpler, or provide better approximation guarantees, or work for more general families of utilities~(see e.g. \citet{SGHSY18}, \citet{ARSW17a}, \citet{AMNS15a}, \citet{BaKr17} and \citet{XiLu19a}).

The second solution concept that has been proposed is a relaxation of envy-freeness called \emph{envy-free up to one item (EF1)}. An allocation satisfies EF1 if it is envy-free or any agent's envy for another agent can be removed if some item is ignored. Under additive utilities, 
EF1 can be achieved by a simple algorithm called the round-robin sequential allocation algorithm. Agents take turn in a round-robin manner and pick their most preferred unallocated item. The interest in EF1 was especially piqued when \citet{CKM+16a} proved that for positive additive utilities, a rule based on maximizing Nash social welfare finds an allocation that is both EF1 and Pareto optimal. The result has led to very interesting followup work. For example, \citet{BKV18a} presented a  pseudo-polynomial time algorithm for the same setting to compute an Pareto optimal and EF1 allocation. The returned allocation also provides a 1.45 approximation of the maximum Nash welfare.\footnote{There is a stream of papers primarily focussed on approximating Nash welfare. For example, \citet{CoGz15a} presented a polynomial-time 2.889-approximation algorithm for maximum Nash welfare.}

A majority of the results on indivisible items assume that the items are goods yielding positive utilities (see e.g. \citet{BCM15a,LaRo16}). 
In recent years, there has been a push to obtain similar existence and algorithmic results for negative utilities (see e.g. \citet{Aziz16a}) or  more general utility functions~\citep{PlRo18,OPS19a}. For example, \citet{ACI+18} designed an algorithm that returns in polynomial time an EF1 allocation for doubly monotonic separable preferences.

There are ongoing efforts to develop models and algorithms that take into account real-life features including distributional constraints, dynamic and online settings,  distributed settings, and asymmetric agents. 
An interesting new setting that has been considered is one in which the agents are partitioned into groups and the items are allocated to groups~\citep{SeSu19a}.
Researchers have also been exploring stronger or alternative fairness and efficiency concepts. For example, there is a renewed focus on concepts that concern envy between groups of agents~\citep{AzRe19a,CFS+19,BCEZ19}. Another direction concerns scenarios where items are viewed as nodes of a graph and only those allocations are considered in which each agent gets a connected coalition~\citep{BCE+17a}.  

\section{Allocation of Divisible Items}
\label{sec:div}

When divisible items are allocated among agents, it is easier to guarantee the existence of fair allocations. It can also be relatively easier to compute a fair allocation. Again, consider the example of two agents and one item.
If the item is divisible, we can simply allocate half of the item to each agent, guaranteeing many reasonable fairness criteria including envy-freeness. 

In contrast to the case of indivisible items, an envy-free and Pareto optimal allocation always exists for the case of divisible items. For positive additive utilities, such an allocation can be computed by maximizing Nash social welfare in polynomial time. When the utilities are additive but negative, there are still important gaps in our understanding of the computational complexity of envy-free and Pareto optimal allocations (see e.g. \citet{BMSY17}). 

For divisible items, one important class of utility functions is called \emph{Leontief preferences} in which each agent requires fixed proportions of the items (such as 1 CPU and 5 GB RAM to perform one task). \citet{GZH+11a} proposed a solution for the problem of allocating as much of each item as possible
while equalizing the agents' use of the fraction of their own dominant (bottleneck) resource. The method satisfies envy-freeness, proportionality, Pareto optimality, and strategyproofness. These solutions have  been extended in important ways such as dealing with indivisibilities and dynamic settings~\citep{PPS12a}.

Another model of divisible items is called cake-cutting~\citep{RoWe98a,Proc15a,Proc12a}. A cake is a unit interval representing a heterogeneous item. 
An agent may have different utilities for different subintervals, even if they are of equal size. The challenge is to compute fair allocations using queried information from agents. Cake-cutting has a very long history dating back to the inception of the mathematical theory of fair division. Two of the most well-studied concepts are envy-freeness and proportionality (each agent gets at least $1/n$ of her value of the whole cake). While there are well-known and simple algorithms to compute a proportional allocation, finding envy-free allocations is more challenging. In the last few years, new algorithms have been proposed that compute an envy-free allocation in a number of queries that is bounded in the number of agents. The algorithms work for either positive utilities~\citep{AzMa16c} or negative utilities~\citep{DFHY18a}. 
There is no known bounded algorithm for the case of mixed utilities. 
It has also been noted by the authors that if the requirement to allocate the whole cake is relaxed, then there are simpler and faster algorithms to compute an allocation that is both envy-free and proportional.  In recent years, several new variants of the cake-cutting problems have been studied (see, e.g \citet{SHA15a}).

Recently, researchers have approached problems where indivisible items are allocated among agents, but a minimal number of items are shared or treated as divisible to guarantee existence of fair allocations~\citep{SaSe19b}. 
The approach can be viewed as a natural extension of the Adjusted Winner rule~(see e.g. \citet{BrTo96a}) that finds an envy-free and Pareto optimal allocation among two agents in which only at most one item is shared. 

Money can also be viewed as a divisible item that can be moved around to achieve fairness. A well-studied discrete allocation problem with money is the fair room-rent division problem (see e.g. \citet{ADG91a,ASU04a,GMPZ17a}).

\section{Two-Sided Matching}
\label{sec:match}

In two sided matching markets~(see e.g. \citet{Manl13a}), the goal is to match entities from two different sets. A typical example of the problem is one where 
students/doctors are matched to school/hospitals. Students have preferences over schools. Schools have priorities over students. An outcome in these settings is a matching of students to schools. 

In two-sided matching, an important criterion of desirable outcomes is a fairness property that requires that no student should want to replace another student in a hospital when the hospital has a higher priority for the former student over the latter student. The concept is also based on the idea of envy. The only difference is that an agent can only envy another agent if the former has higher priority.

The property of fairness along with that of non-wastefulness (no student prefers to take some empty slot in a school) are the 
central properties that underpin much of the work on two-sided matching. 
For many general two-sided matching markets, 
the two properties can be simultaneously satisfied by an allocation returned by the well-known Deferred Acceptance algorithm~\citep{Roth08a} or its generalizations. For several important settings such as the allocation of school seats to students~\citep{AbSo03b}, one version of the algorithm additionally satisfies strategyproofness. 

An important research agenda extends this kind of result to other two-sided matching problems with distributional constraints. These constraints included upper quotas on regions of hospitals~(see e.g. \citet{KaKo15,BFIM10,GIKY+16a}) as well as target lower quotas for diverse representation~\citep{AGSW19a,EHYY14a,KHIY15a}.  These new settings give rise to fundamental questions of tradeoffs between merit, diversity, and segregation. Yet another direction is taking a time-sharing approach to two-sided matching (see e.g. \citet{AzKl19b,KeUn15a}).

\section{Fairness in other Social Choice Settings}
\label{sec:sc}

%The domain of fair allocation is 
We have mostly focussed on allocations problems where agents get ownerships of their allocation. 
We can also consider allocation of public goods such as the funding of public projects. Public goods can be selected via voting on the different options. 
In classical voting settings in which a single alternative is to be selected, it is very difficult to guarantee reasonable representation of sizable minorities. Consider the case where 51$\%$ of the agents most prefer alternative $a$ and 49$\%$ of the agents prefer alternative $b$. Then alternative $a$ that is supported by a majority is selected even though the remaining agents may find it unpalatable. 

There are other voting settings where fairness concerns become more meaningful. One such setting is a simple extension of single-winner voting to selecting multiple winners (see e.g. \citet{ABC+16a,FSST17a,Aziz19b}). Another setting is probabilistic voting or portioning in which each alternative is given a certain probability weight or a fraction of a budget (see e.g. \citet{AAC+19a,ABM19a,Aziz18a,Bran17a}). In both settings, meaningful fairness axioms have been proposed and algorithms have been designed to satisfy the axioms.
A general principle while formalizing the axioms is that of proportional representation: if a large enough group cohesively prefers some outcome(s), then that group should get suitable representation (see e.g. \citet{AzLe19a}).  
Recently, fairness concerns are also being examined in participatory budgeting problems in which projects have costs and there is a constraint on the maximum budget that can be used~\citep{FGM16a,ALT18a,FST+17a}. 

Another collective decision-making setting where fairness becomes more meaningful is the case where there are several issues with each issue having its own candidates outcomes~\citep{CF017}. Agents have preferences over outcomes over each of the issues. The setting is more general than allocation of indivisible items in following way. Each issue corresponds to the decision of allocation of a particular item. The possible outcomes of an issue correspond to giving the item to one of the agents. It turns out that a solution based on maximizing Nash social welfare satisfies both Pareto optimality and a natural relaxation of proportionality. 

\section{Discussion}
\label{sec:discuss}

%I have provided a bird's eye view of multi-agent settings in which fairness is an important issue. 
I discussed some settings where fairness plays an important role. Fairness concerns can come up in many other multi-agent settings as well such as auctions and coalition formation. Much of the work in multi-agent fair allocation is centered around formal axioms to capture fairness.\footnote{In addition to standard axioms such as envy-freeness, a formal theory of fair allocation covers many different axiom for desirable rules or algorithms~(see, e.g. \citet{Thom15a}.)} It will be interesting to use data-centric approaches to understand how humans conceptualize fairness or decide on tradeoffs between societal outcomes. 

Another area of computer science where fairness is taking center stage is machine learning. Machine learning algorithms learn from data to make actionable suggestions such as giving a loan. Whether these algorithm are biased is a prominent concern for society~(see, e.g. \citet{KLR18a}). A natural approach to incorporate fairness in these algorithms follows the Aristotelian ``equal treatment of equals'' principle.
It is an interesting challenge to explore ways in which ideas from multi-agent fair allocation can contribute to fairness in machine learning. In general, the interplay of fairness with other desirable aspects such as accountability, transparency, and trust is broad research question~(see, e.g. \citet{ShPa19a}). 

As computers start to make important decisions in our lives, it is timely to revisit the lessons learnt. New computer applications will give rise to novel and challenging problems so it is critical to continue updating our toolkit of axioms and methods for designing and enabling fair multi-agent systems.

\section*{Acknowledgements}

The survey is based on a presentation given by the author at the Dagstuhl-Seminar on Application-Oriented Computational Social Choice (Seminar Number 19381). The author thanks Alex Lam and the reviewers for useful comments. 

% \clearpage
\bibliographystyle{aaai}
%\bibliographystyle{ACM-Reference-Format}
 % \bibliography{abb,group,aziz,fair,GEF,matching,adt}
 %

\end{document}